Chem1

At 1000 K, iron metal is produced through the reac
Answer #1

the given reaction is

FeO (s) + CO ---> Fe + C02

solids are not considered in the equilibrium constant equation

so

Kc = [C02] / [C0]

now

consider the reaction


FeO + CO ---> Fe + C02

using ICE table

initial conc of CO , C02 are 1 , 0.5

change in conc of CO , C02 are -x , +x

equilibrium conc fo CO , CO2 are 1-x , 0.5 +x

so

we get

Kc = [C02] / [C0]

0.259 = (0.5 +x) / ( 1-x)

0.259 - 0.259x = 0.5 + x

solving we get

x = -0.1914

so

[C0]eq = 1 - x = 1 - (-0.1914) = 1.1914 M


[C02]eq = 0.5 + x = 0.5 - 0.1914 = 0.3086

so

the conc of C02 at equilibrium is 0.309 M

so

the answer is option d) 0.309 M